# Emotional Speech-driven 3D Body Animation via Disentangled Latent Diffusion

### v2

@@ B.4 Training and Losses
[…]
@@
Method EA (%) $\uparrow$ EF1$\uparrow$ SA (%) $\uparrow$ SF1$\uparrow$
Ours 91.531 0.914 96.060 0.960
Emo-disentangle 91.966 0.918
Sty-disentangle 96.095 0.961
[…]
Ours .3739 .3740 .3816 .3815
Emo-disentangle .3793 .3792 .3905
Sty-disentangle .3769 .3770 .3887
Finally, we employ the cross-reconstruction losses for emotion, style, and content. This loss ensures that we can combine any three style, content, and emotion latents and decode them into a valid reconstruction. As shown in Fig. A.1 and Fig. A.2, this is a three part cross reconstruction process. In this process, we extract content $E_{c}(a_{*})$, emotion $E_{e}(a_{*})$, and style $E_{s}(a_{*})$ latents of all four different audios. Given two input audios of the different contents $c_{i}$ and $c_{j}$, with the same speaker, and the same emotion, we swap the emotion latents between the audio pair, and decode the two audios back. Since the emotion class is constant within a quadruple, the emotion cross-reconstruction should be equal to the original audio. Similarly, we cross-reconstruct an input audio with two style latents of the same person, but of different sequence. Enforced by:
@@ B.6 Ablation Experiments
We conduct two ablation studies with the speech disentanglement model. One to only disentangle emotion from content (dropping $\mathcal{L}_{sty}$, $\mathcal{L}_{xsty}$). The other to only disentangle only style from content (dropping $\mathcal{L}_{emo}$, $\mathcal{L}_{xemo}$). Tab. A.1 shows the accuracy and F1 scores for emotion and style latent vectors. The Emotion Accuracy (EA), Style Accuracy (SA), Emotion F1 Score (EF1), and Style F1 Score (SF1) in our speech disentanglement model exhibit only marginal differences compared to the results obtained in the ablation experiments. We report the test set self- and cross-reconstruction errors in Tab. A.2. The cross-reconstruction errors are comparable to self-reconstruction errors which indicates that the individual latents from different audios can be combined to produce valid outputs. This holds for the main model and also the ablated models. However, the ablated models are not able to factor the audio into all three components due to the dropped loss terms. We observe the robust performance of our audio model, by accounting for the complex interplay between emotion and style. By concurrently disentangling three latent vectors, our approach effectively captures the intricate relationships in the audio data, allowing to jointly model and distinguish both emotion and style factors.
@@ C.1 Motion Prior And Latent Denoiser
In this section we include detailed illustrations of the motion prior and latent denoiser. Fig. A.4 illustrates the inference process employed by our model. Fig. A.5 illustrates the forward diffusion and the reverse audio-conditioned denoising process, operating at the latent space. Finally, A.7 shows the diagram of the architecture of the motion prior.
@@ C.2 Methods Trained on Coarse Skeletal Data
We compare AMUSE with methods trained on coarse skeletal data. We choose DSG [103], CaMN [58], Zhu et al. [116] and MoGlow [4] as recent gesture generation models using audio input. AMUSE produces more synchronized gestures and better represents the underlying audio emotion compared to the state of the art methods trained on skeletal data, as shown in our supplementary video. Additionally, these methods are not trained to output 3D meshes. We observe uncanny poses and self-penetrations as shown in Fig. A.8. In our video, we provide additional comparison with these skeleton based methods in both formats, the original predictions of those models and 3D meshes which are created via Inverse Kinematics (IK). We exclude the conversion to 3D mesh for Zhu et al. [116] because the output skeleton format is incompatible with SMPL-X topology.
@@ C.3 Ablation Experiments
. Following the procedure described in the Sec. 5.1 , we report quantitative evaluation scores in Tab. A.3, comparing AMUSE with the ablation models and GT. The version w/o speech disentanglement model produces lower-quality gestures and lacks editing capabilities compared to the complete model. This is because it lacks a component for separating emotion, content and style in the audio. The scores for the ablation models without motion prior are the lowest, indicating that this model did not converge successfully. Additionally, in Tab. A.4 we report improved FGD and Div scores when the motion prior and diffusion model are trained jointly compared to when trained separately, indicating that joint training yields superior results. Furthermore, we conduct additional ablation experiments with and without alignment losses ($\mathcal{L}_{align}$, $\mathcal{L}_{Valign}$). Including alignment losses results in a GA of 46.79%, whereas without them, the GA drops to 30.89%, demonstrating the alignment losses effectiveness. Moreover, we compute the average jerk of the left and right hands for motion sequences belonging to the same audio of [103], ours, and GT, reporting it in m/s3 as 1.18, 1.10, and 0.065, respectively. This signifies that the GT motion is the most steady, whereas ours is slightly smoother over time compared to [103].
@@ Appendix E Gesture Emotions And Semantics
SRGR =$\textup{SRGR }=\lambda\sum\frac{1}{\textup{T x J}}\sum_{t=1}^{T}\sum_{j=1}^{J}\textbf{1}\left[\left\|p_{t}^{j}-\hat{p}_{t}^{j}\right\|_{2}<\delta\right]$
@@
. We obtain the ground-truth semantic score, provided by the BEAT authors, for computing the SRGR. In Fig. A.9, we present semantic scores for the same subject across various emotions, while Fig. A.10 illustrates semantic scores for all subjects expressing the same emotion. This allows us to observe how subjects gesture differently with different emotions and how different subjects gesture for the same emotion. While we acknowledge the high-quality dataset introduced by the BEAT authors, our model has the potential to deliver even better results and improved expressivity with an enhanced dataset quality.
@@
. Gestural emotions are complex, influenced by internal states of subject, social signals, and their perception vary significantly across individuals with diverse cultural backgrounds. AMUSE is designed to capture perceived gestural emotions. While we demonstrate AMUSE with a gesture emotion recognition accuracy of 46.76% and AMUSE-Edit with 34.18%, outperforming other state of the art methods, it is important to note that recognizing emotion from gestures remains a challenging task in computer vision. We observe the gesture emotion accuracy for the GT sequence is 64.04%. There is still ample room for improvement in addressing this complex problem. Additionally, in Fig. A.12, we present the confusion matrix for ground truth (GT) emotion predictions on the left and reconstructions (gestures generated using the original style, emotion, and content latents of a given audio) on the right. We observe a robust correlation between the predictions on GT and the reconstructions for all eight emotions. Additionally, we conducted experiments on gesture edits by swapping emotion latents from one audio with those from another audio of the same subject but with a different utterance. In Fig. A.11, we showcase two exemplars, transforming from Happy to rest and Surprise to rest. Given the diversity of eight emotions, gestural edits offer numerous possibilities, rendering this a broad and challenging problem. Although Fig. A.11 displays promising results for emotion label predictions with clear diagonal pattern of the confusion matrix, we acknowledge the inherent difficulty in solving this problem.
@@ Appendix F Data Preparation
Method\;$\downarrow$ Ours Others | Ours Others
@@
Appendix G Review of State of the Art Methods
Data selection and input formats.
@@ Appendix H Additional Perceptual Study Details
Here we describe additional details of the AMT study reported in the Sec. 5.3. We show aggregate preference scores in Tab. A.5. AMUSE outperforms all methods compared against in both criteria - synchronization with the speech and the appropriateness with respect to the specified emotion. In contrast, Ground Truth (GT) consistently outperforms AMUSE in both tasks. This outcome emphasizes the complexity of the problem, where achieving synchrony with speech and meeting specified emotional appropriateness remain challenging objectives.

### v3

@@ B.4 Training and Losses
$\begin{aligned} \mathcal{L}_{emo}&=-\mkern-18.0mu\sum_{\begin{subarray}{c}1\leq l_{e}\leq n_{e}\end{subarray}}\mkern-18.0muy_{l_{e}}\log(p_{l_{e}}),\\ \mkern 9.0mu\mathcal{L}_{sty}&=-\mkern-18.0mu\sum_{\begin{subarray}{c}1\leq l_{s}\leq n_{s}\end{subarray}}\mkern-18.0muy_{l_{s}}\log(p_{l_{s}}),\end{aligned}$
@@
Finally, we employ the cross-reconstruction losses for emotion, style, and content. This loss ensures that we can combine any three style, content, and emotion latents and decode them into a valid reconstruction. As shown in Fig.˜A.1 and Fig.˜A.2, this is a three part cross reconstruction process. In this process, we extract content $E_{c}(a_{*})$, emotion $E_{e}(a_{*})$, and style $E_{s}(a_{*})$ latents of all four different audios. Given two input audios of the different contents $c_{i}$ and $c_{j}$, with the same speaker, and the same emotion, we swap the emotion latents between the audio pair, and decode the two audios back. Since the emotion class is constant within a quadruple, the emotion cross-reconstruction should be equal to the original audio. Similarly, we cross-reconstruct an input audio with two style latents of the same person, but of different sequence. Enforced by:
@@ B.5 Implementation Details
[…]
Ours 91.531 0.914 96.060 0.960
Emo-disentangle 91.966 0.918
Sty-disentangle 96.095 0.961
[…]
Ours .3739 .3740 .3816 .3815
Emo-disentangle .3793 .3792 .3905
Sty-disentangle .3769 .3770 .3887
@@ B.6 Ablation Experiments
We conduct two ablation studies with the speech disentanglement model. One to only disentangle emotion from content (dropping $\mathcal{L}_{sty}$, $\mathcal{L}_{xsty}$). The other to only disentangle only style from content (dropping $\mathcal{L}_{emo}$, $\mathcal{L}_{xemo}$). Tab.˜A.1 shows the accuracy and F1 scores for emotion and style latent vectors. The Emotion Accuracy (EA), Style Accuracy (SA), Emotion F1 Score (EF1), and Style F1 Score (SF1) in our speech disentanglement model exhibit only marginal differences compared to the results obtained in the ablation experiments. We report the test set self- and cross-reconstruction errors in Tab.˜A.2. The cross-reconstruction errors are comparable to self-reconstruction errors which indicates that the individual latents from different audios can be combined to produce valid outputs. This holds for the main model and also the ablated models. However, the ablated models are not able to factor the audio into all three components due to the dropped loss terms. We observe the robust performance of our audio model, by accounting for the complex interplay between emotion and style. By concurrently disentangling three latent vectors, our approach effectively captures the intricate relationships in the audio data, allowing to jointly model and distinguish both emotion and style factors.
@@ C.1 Motion Prior And Latent Denoiser
In this section we include detailed illustrations of the motion prior and latent denoiser. Fig.˜A.4 illustrates the inference process employed by our model. Fig.˜A.5 illustrates the forward diffusion and the reverse audio-conditioned denoising process, operating at the latent space. Finally, A.7 shows the diagram of the architecture of the motion prior.
@@ C.2 Methods Trained on Coarse Skeletal Data
We compare AMUSE with methods trained on coarse skeletal data. We choose DSG [103], CaMN [58], Zhu et al. [116] and MoGlow [4] as recent gesture generation models using audio input. AMUSE produces more synchronized gestures and better represents the underlying audio emotion compared to the state of the art methods trained on skeletal data, as shown in our supplementary video. Additionally, these methods are not trained to output 3D meshes. We observe uncanny poses and self-penetrations as shown in Fig.˜A.8. In our video, we provide additional comparison with these skeleton based methods in both formats, the original predictions of those models and 3D meshes which are created via Inverse Kinematics (IK). We exclude the conversion to 3D mesh for Zhu et al. [116] because the output skeleton format is incompatible with SMPL-X topology.
@@ C.3 Ablation Experiments
. Following the procedure described in the Sec. 5.1 , we report quantitative evaluation scores in Tab.˜A.3, comparing AMUSE with the ablation models and GT. The version w/o speech disentanglement model produces lower-quality gestures and lacks editing capabilities compared to the complete model. This is because it lacks a component for separating emotion, content and style in the audio. The scores for the ablation models without motion prior are the lowest, indicating that this model did not converge successfully. Additionally, in Tab.˜A.4 we report improved FGD and Div scores when the motion prior and diffusion model are trained jointly compared to when trained separately, indicating that joint training yields superior results. Furthermore, we conduct additional ablation experiments with and without alignment losses ($\mathcal{L}_{align}$, $\mathcal{L}_{Valign}$). Including alignment losses results in a GA of 46.79%, whereas without them, the GA drops to 30.89%, demonstrating the alignment losses effectiveness. Moreover, we compute the average jerk of the left and right hands for motion sequences belonging to the same audio of [103], ours, and GT, reporting it in m/s3 as 1.18, 1.10, and 0.065, respectively. This signifies that the GT motion is the most steady, whereas ours is slightly smoother over time compared to [103].
@@ Appendix E Gesture Emotions And Semantics
$\textup{SRGR }=\lambda\sum\frac{1}{\textup{T x J}}\sum_{t=1}^{T}\sum_{j=1}^{J}\textbf{1}\left[\left\|p_{t}^{j}-\hat{p}_{t}^{j}\right\|_{2}<\delta\right]$
@@
. We obtain the ground-truth semantic score, provided by the BEAT authors, for computing the SRGR. In Fig.˜A.9, we present semantic scores for the same subject across various emotions, while Fig.˜A.10 illustrates semantic scores for all subjects expressing the same emotion. This allows us to observe how subjects gesture differently with different emotions and how different subjects gesture for the same emotion. While we acknowledge the high-quality dataset introduced by the BEAT authors, our model has the potential to deliver even better results and improved expressivity with an enhanced dataset quality.
@@
. Gestural emotions are complex, influenced by internal states of subject, social signals, and their perception vary significantly across individuals with diverse cultural backgrounds. AMUSE is designed to capture perceived gestural emotions. While we demonstrate AMUSE with a gesture emotion recognition accuracy of 46.76% and AMUSE-Edit with 34.18%, outperforming other state of the art methods, it is important to note that recognizing emotion from gestures remains a challenging task in computer vision. We observe the gesture emotion accuracy for the GT sequence is 64.04%. There is still ample room for improvement in addressing this complex problem. Additionally, in Fig.˜A.11, we present the confusion matrix for ground truth (GT) emotion predictions on the left and reconstructions (gestures generated using the original style, emotion, and content latents of a given audio) on the right. We observe a robust correlation between the predictions on GT and the reconstructions for all eight emotions. Additionally, we conducted experiments on gesture edits by swapping emotion latents from one audio with those from another audio of the same subject but with a different utterance. In Fig.˜A.12, we showcase two exemplars, transforming from Happy to rest and Surprise to rest. Given the diversity of eight emotions, gestural edits offer numerous possibilities, rendering this a broad and challenging problem. Although Fig.˜A.12 displays promising results for emotion label predictions with clear diagonal pattern of the confusion matrix, we acknowledge the inherent difficulty in solving this problem.
@@ Appendix F Data Preparation
Appendix G Review of State of the Art Methods
Data selection and input formats.
@@ Appendix G Review of State of the Art Methods
[…]
@@ Appendix H Additional Perceptual Study Details
Here we describe additional details of the AMT study reported in the Sec. 5.3. We show aggregate preference scores in Tab.˜A.5. AMUSE outperforms all methods compared against in both criteria - synchronization with the speech and the appropriateness with respect to the specified emotion. In contrast, Ground Truth (GT) consistently outperforms AMUSE in both tasks. This outcome emphasizes the complexity of the problem, where achieving synchrony with speech and meeting specified emotional appropriateness remain challenging objectives.